Came across a problem with fractions. The problem is 6/5(z)=4/7, what would z equal.
Answer by stanbon(75887)

6/5(z)=4/7, what would z equal
----------------
Multiply both sides by (5/6) to get:
z = (5/6)(4/7)
---
z = 20/42
---
z = 10/21
===============
Cheers,
Stan H.
=============